OpenStudy (anonymous):

A particle is moving with velocity v(t) = t2 – 9t + 18 with distance, s measured in meters, left or right of zero, and t measured in seconds, with t between 0 and 8 seconds inclusive. The position at time t = 0 sec is 1 meter right of zero, that is, s(0) = 1. The average velocity over the interval 0 to 8 seconds The instantaneous velocity and speed at time 5 secs The time interval(s) when the particle is moving right The time interval(s) when the particle is going faster slowing down Find the total distance the particle has traveled between 0 and 8 seconds

OpenStudy (freckles):

Do you need help with all of these?

OpenStudy (freckles):

well the instantaneous velocity at t=5 can be found by pluggin in t=5 the velocity function is the one given

OpenStudy (anonymous):

yes please

OpenStudy (freckles):

so you found the instantaneous velocity yet?

OpenStudy (anonymous):

yes which is 88

OpenStudy (freckles):

isn't your equation v=t^2-9t+18

OpenStudy (freckles):

I don't get 88 when pluggin in 5

OpenStudy (anonymous):

i rechecked is -2

OpenStudy (freckles):

sounds tons better :

OpenStudy (freckles):

speed=|v|

OpenStudy (anonymous):

so is 2

OpenStudy (freckles):

so yeah your second question is actually a whole bunch easier than your first question

OpenStudy (freckles):

because to find the average velocity we need to find the position function

OpenStudy (freckles):

this is because average velocity on [0,8] equals: \[\frac{s(8)-s(0)}{8-0}\]

OpenStudy (freckles):

the velocity normally would have been found by finding the derivative of the position function s but that was given to us so it made our work for instantaneous rate of change and speed tons easy

OpenStudy (freckles):

do you know how to find s given s'?

OpenStudy (anonymous):

2t-t ?

OpenStudy (freckles):

if it wasn't clear from what i was saying i'm saying velocity=position' v=s' so to find s you need to integrate both sides .

OpenStudy (freckles):

\[s'=v \\ \int\limits s' dt=\int\limits v dt \\ s(t)=\int\limits v dt \]

OpenStudy (anonymous):

freckles [Math Processing Error]

OpenStudy (freckles):

have you integrated v yet ?

OpenStudy (anonymous):

2t-t? i don't know how to do it !:(

OpenStudy (anonymous):

2t-9

OpenStudy (anonymous):

8t^2-72t+144 ?

OpenStudy (freckles):

do you have any experience with power rule ?

OpenStudy (freckles):

\[\int\limits t^n dt=\frac{t^{n+1}}{n+1}+C , n \neq -1 \]

OpenStudy (freckles):

the first term you have t^2 use the above formula

OpenStudy (freckles):

you can also use it for the second term -9t

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3t^2/3 +c ?

OpenStudy (freckles):

2+1 is 3 not 2 and the 3 in front shouldn't magically appear

OpenStudy (freckles):

\[\int\limits_{}^{}t^2 dt=\frac{t^{2+1}}{2+1}+C=\frac{t^3}{3}+C\]

OpenStudy (anonymous):

i thought i have to solve for derivetive

OpenStudy (freckles):

no like s'=v so the s=antiderivative of v

OpenStudy (freckles):

@satellite73 I'm going to go if you want to help this person in calculus. I know how much you like calculus. (It is just velocity and speed and all that jazz, nothing too scary).